General Physics

raindrops of radius 1 mm and mass 4 mg are falling with a speed of 30 m/s on head of a bald person.The drops splash on the head and come to rest.assuming equivalently that the drops cover a distance equal to their on the head,estimate the force excerted by each drop on head.

Aditya
12 Years agoGrade

1 Answer

Binay kumar singh
9 Years ago
Radius = 1mm = 10^-3 m; m = 4mg= 4×10^-6; u= 30m/s; v =0. V^2- u^2 = 2as or a = -450× 10^3m/s^2 retardation.F= ma = 4×10^-6× 450× 10^3 = 1.8N